# Normalizing Flow-Assisted Nested Sampling on Type-II Seesaw Model

### v1

@@ IV.2.2 Yukawa matrix from neutrino oscillation data
We have only discussed the scalar sector up to this point. Once the triplet VEV in known, one can determine the neutrino Yukawa matrix using Eq. 12 such that the light neutrino masses and mixing angles are consistent with the neutrino oscillation data [40]. We have performed the calculation for both normal (NO) and inverted ordering (IO) of light neutrino masses. The neutrino oscillation data is made up of two mass-square differences, three mixing angles, and a CP-violating phase angle ($\delta$).
@@
Eq. 13 shows the resultant neutrino Yukwa matrix elements, subjected to neutrino oscillation data and for the choice of $v_{T}\sim 0.5-2.5$ GeV favored by the constraints at $95\%$ CL from our analysis.
@@
We have kept $\delta=0$ in this calculation and have assumed NO in neutrino masses. The hierarchy is clearly reflected in the resultant Yukawa elements. The central tendencies (mean) and associated dispersions (standard deviation, $\sigma$) corresponding to all elements are quoted. As expected, given the favored region of $v_{T}$, the Yukawa elements are all $\sim 10^{-10}$ with even smaller off-diagonal entries. This ensures that all LNV or LFV decays are well suppressed. Note that, one can perform a similar calculation with a non-zero $\delta/\pi=194^{+24}_{-22}$ [40]. The phase angle will require the Yukawa matrix to be complex. However, the large uncertainty associated with it results in large, flat posteriors, with overlapping $68\%$ and $95\%$ regions. We show posterior distributions of the Yukawa elements for the NO scenario with $\delta=0$ in Fig. 9. One can generate similar corner plots for the IO scenario and with non-zero $\delta$. These results, along with all other results (plots, samples, and trained networks), are available in the GitHub repository.
@@
IV.2.3 Sensitivity of the Observables to the parameters
@@
In Fig. 8 we show the sensitivity of some precisely measured observables to the obtained parameter space. To get the distribution of the observables in these plots, the whole posterior—allowed by all theoretical and experimental constraints—is used and a Bi-normal distribution is estimated. On the left, we show the overlap of the observable distribution (the red ellipse showing $68\%$ credible interval) calculated from the posterior with the $68\%$ and $95\%$ experimental ranges of $\rho$ parameter and the SM-like Higgs mass $m_{h}$. On the right, we show the equivalent region on the $S-T$ plane (red ellipse) and compare it with the $68\%$ CL computed by the GFitter group [86] and it is clear that the high likelihood region obtained through our analysis is well within the experimental electroweak precision bounds. Regarding the $\rho$-parameter, we started our analysis with the estimate made by the Particle Data Group (PDG) in 2022 which was $\rho=1.00039\pm 0.00019$ [87]. While we were performing our analysis, PDG updated their estimate which now is quoted as $\rho=1.00031\pm 0.00019$ [88]. It does not affect our analysis or results. In fact, as the figure shows, the new measurement is in even better agreement with our estimated parameter space.
@@
To check the sensitivity of our analysis to the parameters, we take the help of the ‘SHapley Additive exPlanations’ (SHAP) [133]. It is a widely used method to calculate the relative importance of features (parameters, for our case) in obtaining a prediction (observables/likelihood for us), using game-theoretic approaches. To use this in our analysis, we first use our final sample from posterior (allowed by HiggsBounds, $>93000$ points) to obtain the respective negative log-likelihoods, which together make both training and test sets. After successfully training a Gradient Boosted Trees algorithm (we could have used some sophisticated network as well, but the SHAP would hardly change), we apply it to the test dataset to obtain the predictions and corresponding measurements. The ML algorithm is just a placeholder here for a non-linear predictor function. We do the SHAP analysis with these.
@@
A vector of Shapley values is calculated for every point prediction, with one component for each of the features. These values signify the relative importance of the respective feature in obtaining that prediction using both the impact of the change in that feature on the output, and the distribution of that feature. With one of the features fixed at a specific value, the corresponding Shapley value is the difference between the expected prediction (mean prediction for all features varied) and the mean of all predictions with that feature fixed at the value mentioned before (partial dependence). In game-theoretic parlance, Shapley values, when added up, always provide the difference between the game outcome with all players present and with no players present (the difference between the actual prediction and the mean prediction). This is demonstrated in the top chart (waterfall plot) of Fig. 10. The point prediction considered is one of the highest likelihood points in our obtained NS dataset. The lower panel of Fig. 10 shows the distributions of all the Shapley values for each parameter with their relative impact on the $x$-axis. The feature values are shown with a gradient color code.
@@ V Summary and Conclusion
Sampling a high-dimensional parameter space, as is often encountered with various BSM scenarios, is a daunting task. The most widely used statistical tool to perform this and draw Bayesian inference is MCMC. In spite of its effectiveness in performing the task, it remains computationally expensive to find all the modes when dealing with a multi-modal parameter space. Parallelization and automatic tuning of the convergence are also non-trivial. NS is a more efficient technique that can be used to alleviate these issues. NS creates a weighted sample for the posterior starting from the whole constrained prior and is better suited to handle multi-modal problems. Its convergence is automatically tunable and it is highly parallelizable. The algorithm, however, slows down increasingly as it approaches high-likelihood regions, thus increasing the time to converge. A rapidly decreasing volume of the live region compared to the constrained prior is one of the main reasons of this slowdown. The main source of slowdown in the case of BSM searches, however, is the HEP spectrum generators. Though these need to be run at each parameter point, a large chunk of any BSM parameter space gives rise to un-physical particle spectra and is therefore discarded. To reduce unnecessary runs of the spectrum generator, we propose using an ML classifier trained to predict whether a random set of input parameters can give rise to a theoretically consistent particle spectrum. We use an ensemble SNN, obtained from a chosen structure after hyper-parameter tuning, for the classifier. In addition to this, we introduce a generative algorithm within the NS loop to iteratively learn the structure of the coupled space of parameters and the log-likelihood function. This produces a large number of high-likelihood candidates with increasing accuracy and calculating the actual likelihood on these prevents the slowdown as the algorithm approaches convergence. A special type of normalizing flow network, known as Real-NVP or RNVP, is used to as the generative algorithm. This network is trained repetitively with truth data which contains the parameters and their actual likelihoods. This is trained not at each iteration, but only after the pooled dataset has enough samples saved. The convergence depends crucially on the way the training dataset is chosen from the pooled data. The method labeled RNVP-SF optimizes the runtime while generating the correct posterior.
@@
We have considered the Type-II Seesaw model as a representative BSM scenario. This is a well-motivated and widely studied BSM scenario from the viewpoint of neutrino mass generation, LFV and LNV decays, and a rich scalar sector. We only consider a relatively large triplet VEV ($10^{-4}$ $3.5$ GeV) scenario that results in a suppression of LFV and LNV decays. Our focus has exclusively been on the posterior of scalar sector parameters, subjected to the 125 GeV Higgs data, $\rho$-parameter, and the oblique parameters. We present our results in terms of $2D$ marginal distributions on various parameter planes and show the impact of direct collider search results alongside them. We have also shown the preferred mass regions for the exotic scalars, allowed by all theoretical and experimental constraints. We observe that for all these allowed points up to $95\%$ CL, the exotic scalar masses lie above $\sim 350$ GeV with most of them clustered in the $\sim 400-800$ GeV range. The mass difference between the doubly and singly charged Higgs mass lies within $\lesssim 40$ GeV window on either side of zero. The mixing angle between the two CP-even Higgs for the $95\%$ allowed region of $v_{T}$ ($\sim 0.3-2.5$ GeV) is small (as expected) with $\sin\alpha\sim 10^{-3}$. We further check the updated collider constraints on doubly charged Higgs masses and observe that the most stringent constraint arises when they are pair-produced and decay into a pair of $W$-bosons. However, the favored parameter region obtained here is yet to be probed at the existing luminosity. We also comment on how the favored parameter space affects the choice of neutrino Yukawa matrix elements when confronted with the neutrino oscillation data with a zero and non-zero CP-violating phase.
@@
The sampling technique described here is only the first iteration of its kind and it has the potential to improve in various ways. On one hand, the generative algorithm can be improved with better networks (e.g. Masked Auto-regressive Flows [134]) in favor of more precise predictions; on the other hand, the simple region sampler can be upgraded (e.g. with multi-ellipsoid search algorithms, such as MultiNest [27]). Even in its infancy, the method is unique and extremely efficient, as is evident from the results. It was run in a single Windows desktop, with SPheno called through a virtual UNIX environment (WSL). With a complete Linux implementation, it was found to be at least 5 times faster. This reduction in computational time and cost helps the user identify the model posterior easily, thus ensuring a precise prediction of the observables. As a result, the experimental collaborations can concentrate further on the favored parameter space only rather than constructing simplified scenarios representing different parts of the parameter space. With the accumulation of more data, the parameter space can be restricted more effectively, resulting in even more precise predictions. Eventually, in the absence of any new physics signal, this provides an efficient way to discard a model completely. On the other hand, using existing data, one can use this framework to compare similar models and identify the most favored candidate.
Acknowledgements.
@@
RB, SM, and SKP would like to acknowledge support from ANRF (erstwhile DST-SERB), India (grant order no. CRG/2022/003208). SM would like to acknowledge ANRF also for grant order no. CRG/2023/008570. RB and SR acknowledge the invaluable insights and guidance gained during Machine Learning for Particle and Astroparticle Physics (ML4HEP-2024) Workshop, at the Institute of Physics, Bhubaneswar.

### v2

@@ 4.2.2 Yukawa matrix from neutrino oscillation data
We have only discussed the scalar sector up to this point. Once the triplet VEV in known, one can determine the neutrino Yukawa matrix using Eq. 12 such that the light neutrino masses and mixing angles are consistent with the neutrino oscillation data deSalas:2020pgw . We have performed the calculation for both normal (NO) and inverted ordering (IO) of light neutrino masses. The neutrino oscillation data is made up of two mass-square differences, three mixing angles, and a CP-violating phase angle ($\delta$).
@@
Eq. 13 shows the resultant neutrino Yukwa matrix elements, subjected to neutrino oscillation data and for the choice of $v_{T}\sim 0.5-2.5$ GeV favoured by the constraints at $95\%$ CL from our analysis.
@@
We have kept $\delta=0$ in this calculation and have assumed NO in neutrino masses. The hierarchy is clearly reflected in the resultant Yukawa elements. The central tendencies (mean) and associated dispersions (standard deviation, $\sigma$) corresponding to all elements are quoted. As expected, given the favoured region of $v_{T}$, the Yukawa elements are all $\sim 10^{-10}$ with even smaller off-diagonal entries. This ensures that all LNV or LFV decays are well suppressed. Note that, one can perform a similar calculation with a non-zero $\delta/\pi=194^{+24}_{-22}$ deSalas:2020pgw . The phase angle will require the Yukawa matrix to be complex. However, the large uncertainty associated with it results in large, flat posteriors, with overlapping $68\%$ and $95\%$ regions. We show posterior distributions of the Yukawa elements for the NO scenario with $\delta=0$ in Fig. 9. One can generate similar corner plots for the IO scenario and with non-zero $\delta$. These results, along with all other results (plots, samples, and trained networks), are available in the GitHub repository.
@@
4.2.3 Sensitivity of the Observables to the parameters
@@
In Fig. 8 we show the sensitivity of some precisely measured observables to the obtained parameter space. To get the distribution of the observables in these plots, the whole posterior—allowed by all theoretical and experimental constraints—is used and a Bi-normal distribution is estimated. On the left, we show the overlap of the observable distribution (the red ellipse showing $68\%$ credible interval) calculated from the posterior with the $68\%$ and $95\%$ experimental ranges of $\rho$ parameter and the SM-like Higgs mass $m_{h}$. On the right, we show the equivalent region on the $S-T$ plane (red ellipse) and compare it with the $68\%$ CL computed by the GFitter group Haller:2018nnx and it is clear that the high likelihood region obtained through our analysis is well within the experimental electroweak precision bounds. Regarding the $\rho$-parameter, we started our analysis with the estimate made by the Particle Data Group (PDG) in 2022 which was $\rho=1.00039\pm 0.00019$ ParticleDataGroup:2022pth . While we were performing our analysis, PDG updated their estimate which now is quoted as $\rho=1.00031\pm 0.00019$ ParticleDataGroup:2024cfk . It does not affect our analysis or results. In fact, as the figure shows, the new measurement is in even better agreement with our estimated parameter space.
@@
To check the sensitivity of our analysis to the parameters, we take the help of the ‘SHapley Additive exPlanations’ (SHAP) SHAP1 . It is a widely used method to calculate the relative importance of features (parameters, for our case) in obtaining a prediction (observables/likelihood for us), using game-theoretic approaches. To use this in our analysis, we first use our final sample from posterior (allowed by HiggsBounds, $>93000$ points) to obtain the respective negative log-likelihoods, which together make both training and test sets. After successfully training a Gradient Boosted Trees algorithm (we could have used some sophisticated network as well, but the SHAP would hardly change), we apply it to the test dataset to obtain the predictions and corresponding measurements. The ML algorithm is just a placeholder here for a non-linear predictor function. We do the SHAP analysis with these.
@@
A vector of Shapley values is calculated for every point prediction, with one component for each of the features. These values signify the relative importance of the respective feature in obtaining that prediction using both the impact of the change in that feature on the output, and the distribution of that feature. With one of the features fixed at a specific value, the corresponding Shapley value is the difference between the expected prediction (mean prediction for all features varied) and the mean of all predictions with that feature fixed at the value mentioned before (partial dependence). In game-theoretic parlance, Shapley values, when added up, always provide the difference between the game outcome with all players present and with no players present (the difference between the actual prediction and the mean prediction). This is demonstrated in the top chart (waterfall plot) of Fig. 10. The point prediction considered is one of the highest likelihood points in our obtained NS dataset. The lower panel of Fig. 10 shows the distributions of all the Shapley values for each parameter with their relative impact on the $x$-axis. The feature values are shown with a gradient colour code.
@@ 5 Summary and Conclusion
Sampling a high-dimensional parameter space, as is often encountered with various BSM scenarios, is a daunting task. The most widely used statistical tool to perform this and draw Bayesian inference is MCMC. Despite its effectiveness in performing the task, it remains computationally expensive to find all the modes when dealing with a multi-modal parameter space. Parallelization and automatic tuning of the convergence are also non-trivial. NS is a more efficient technique that can be used to alleviate these issues. NS creates a weighted sample for the posterior starting from the whole constrained prior and is better suited to handle multi-modal problems. Its convergence is automatically tunable and it is highly parallelizable. The algorithm, however, slows down increasingly as it approaches high-likelihood regions, thus increasing the time to converge. A rapidly decreasing volume of the live region compared to the constrained prior is one of the main reasons for this slowdown. The main source of slowdown in the case of BSM searches, however, is the HEP spectrum generators. Though these need to be run at each parameter point, a large chunk of any BSM parameter space gives rise to un-physical particle spectra and is therefore discarded. To reduce unnecessary runs of the spectrum generator, we propose using an ML classifier trained to predict whether a random set of input parameters can give rise to a theoretically consistent particle spectrum. We use an ensemble SNN, obtained from a chosen structure after hyper-parameter tuning, for the classifier. In addition to this, we introduce a generative algorithm within the NS loop to iteratively learn the structure of the coupled space of parameters and the log-likelihood function. This produces a large number of high-likelihood candidates with increasing accuracy and calculating the actual likelihood on these prevents the slowdown as the algorithm approaches convergence. A special type of normalizing flow network, known as Real-NVP or RNVP, is used as the generative algorithm. This network is trained repetitively with truth data which contains the parameters and their actual likelihoods. This is trained not at each iteration, but only after the pooled dataset has enough samples saved. The convergence depends crucially on the way the training dataset is chosen from the pooled data. The method labelled RNVP-SF optimizes the runtime while generating the correct posterior.
@@
We have considered the Type-II Seesaw model as a representative BSM scenario. This is a well-motivated and widely studied BSM scenario from the viewpoint of neutrino mass generation, LFV and LNV decays, and a rich scalar sector. We only consider a relatively large triplet VEV ($10^{-4}$ $3.5$ GeV) scenario that results in a suppression of LFV and LNV decays. Our focus has exclusively been on the posterior of scalar sector parameters, subjected to the 125 GeV Higgs data, $\rho$-parameter, and the oblique parameters. We present our results in terms of $2D$ marginal distributions on various parameter planes and show the impact of direct collider search results alongside them. We have also shown the preferred mass regions for the exotic scalars, allowed by all theoretical and experimental constraints. We observe that for all these allowed points up to $95\%$ CL, the exotic scalar masses lie above $\sim 350$ GeV with most of them clustered in the $\sim 400-800$ GeV range. The mass difference between the doubly and singly charged Higgs mass lies within $\lesssim 40$ GeV window on either side of zero. The mixing angle between the two CP-even Higgs for the $95\%$ allowed region of $v_{T}$ ($\sim 0.3-2.5$ GeV) is small (as expected) with $\sin\alpha\sim 10^{-3}$. We further check the updated collider constraints on doubly charged Higgs masses and observe that the most stringent constraint arises when they are pair-produced and decay into a pair of $W$-bosons. However, the favoured parameter region obtained here is yet to be probed at the existing luminosity. We also comment on how the favoured parameter space affects the choice of neutrino Yukawa matrix elements when confronted with the neutrino oscillation data with a zero and non-zero CP-violating phase.
@@
The sampling technique described here is only the first iteration of its kind and it has the potential to improve in various ways. On one hand, the generative algorithm can be improved with better networks (e.g. Masked Auto-regressive Flows papamakarios2017masked ) in favour of more precise predictions; on the other hand, the simple region sampler can be upgraded (e.g. with multi-ellipsoid search algorithms, such as MultiNest Feroz:2008xx ). Even in its infancy, the method is unique and extremely efficient, as is evident from the results. It was run in a single Windows desktop, with SPheno called through a virtual UNIX environment (WSL). With a complete Linux implementation, it was found to be at least 5 times faster. This reduction in computational time and cost helps the user identify the model posterior easily, thus ensuring a precise prediction of the observables. As a result, the experimental collaborations can concentrate further on the favoured parameter space only rather than constructing simplified scenarios representing different parts of the parameter space. With the accumulation of more data, the parameter space can be restricted more effectively, resulting in even more precise predictions. Eventually, in the absence of any new physics signal, this provides an efficient way to discard a model completely. On the other hand, using existing data, one can use this framework to compare similar models and identify the most favoured candidate.
Acknowledgment
@@
RB, SM, and SKP would like to acknowledge support from ANRF (erstwhile DST-SERB), India (grant order no. CRG/2022/003208). SM would also like to acknowledge ANRF for grant order no. CRG/2023/008570. RB and SR acknowledge the invaluable insights and guidance gained during the Machine Learning for Particle and Astroparticle Physics (ML4HEP-2024) Workshop, at the Institute of Physics, Bhubaneswar. The authors also want to thank Jyotishka Datta of Virginia Tech. for explaining to them the nuances of Nested Sampling.